Current Position Analysis

The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $627,000 $494,000
  Marketable securities 726,000 555,800
  Accounts and notes receivable (net) 297,000 185,200
  Inventories 544,500 277,500
  Prepaid expenses 280,500 177,500
  Total current assets $2,475,000 $1,690,000
Current liabilities:
  Accounts and notes payable
  (short-term) $435,000 $455,000
  Accrued liabilities 315,000 195,000
  Total current liabilities $750,000 $650,000

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $ $
2. Current ratio
3. Quick ratio
Answer #1

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $2475000-750000 = 1725000 $1690000-650000 = 1040000
2. Current ratio 2475000/750000 = 3.3 1690000/650000 = 2.6
3. Quick ratio 1650000/750000 = 2.2 1235000/650000 = 1.9